If possible, please show work. I am trying to understand how to get to the correct answers but keep getting it wrong. :(

Consider the following situation: Nicole is a financial analyst in Bidget Corp. As part of her analysis of the annual distribution policy and its impact on the firms value, she makes the following calculations and observations: The company generated a free cash flow (FCF) of $75 million in its most recent fiscal year · The firms cost of capital (WACC) is 15%. The firm has been growing at 10% for the past six years but is expected to grow at a constant rate of 8% in the future . The firm has 18.75 million shares outstanding . The company has $200 million in debt and $125 million in preferred stock Along with the rest of the finance team, Nicole has been part of board meetings and knows that the company is planning to distribute $120 million, which is invested in short-term investments, to its shareholders by buying back stock from its shareholders. Nicole also observed that, at this point, apart from the $120 million in short-term investments, the firm has no other nonoperating assets Using results from Nicoles calculations and observations, solve for the values in the following tables. Select the best answer provided in the selection list. Value Value Value of the firms operations Intrinsic value of equity immediately prior to stock repurchase Intrinsic stock price immediately prior to the stock repurchase Number of shares repurchased Intrinsic value of equity immediately after the stock repurchase Intrinsic stock price immediately after the stock repurchase Based on your understanding of stock repurchases, identify whether the following statement is true or false: If a firm pays a dividend of $0.59 per share, the firms stock price will also fall by $0.59 per share This statement is because if a firm pays a dividend of $0.59 per share, the price per share of the firms stock will also fall by $0.59 to any arbitrage opportunities
Answer #1

Value of firm's operations = FCF0 x (1 + g) / (r - g) = 75 x (1 + 8%) / (15% - 8%) = 1,157.14 million

Value of equity = Value of operations + Cash - Debt - Preferred Stock = 1157.14 + 120 - 200 - 125 = 952.14 million

Stock Price = Value of equity / No. of shares = 952.14 / 18.75 = $50.78

No. of shares repurchased = 120 / 50.78 = 2.36 million shares

Value of equity = Value of operations - Debt - Preferred Stock = 1157.14 - 200 - 125 = 832.14

Stock Price = 832.14 / (18.75 - 2.36) = $50.78

The statement is correct because the stock will fall to prevent any arbitrage opportunities.